Platt Company produces one product, a putter called PAR-putter. Platt uses a standard cost system and determines that it should take one hour of direct labor to produce one PAR-putter. Th The total budgeted overhead at normal capacity is $500,000 comprised of $200,000 of variable costs and $300,000 of fixed costs. Platt applies overhead on the basis of direct labor hours. e normal production capacity for this putter is 100,000 units per year. During the current year, Platt produced 85,000 putters, worked 89,000 direct labor hours, and incurred variable overhead costs of $160,000 and fixed overhead costs of $300,000. Compute the predetermined variable overhead rate and the predetermined fixed overhead rate. Variable Overhead Rate Fixed Overhead Rate Compute the applied overhead for Platt for the year. Applied Overhead Compute the total overhead variance Total Overhead Variance Click if you would like to Show Work for this question: Open Show Work
Answer #1

Predetermine overhead rate :

Variable overhead rate = 200000/100000 = 2 per labour hour

Fixed overhead rate = 300000/100000 = 3 per labour hour

Applied overhead = 85000*1*5 = 425000

Total overhead variance = 425000-460000 = 35000 Unfavorable